OpenStudy (anonymous):

The revenue R in dollars received from selling x radios is R(x)=x(90-x). What number of radios sold will maximize revenue? What is the maximum revenue?

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

multiply this out and get \[R(x)=90x-x^2\] then the minimum will be at the vertex first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{90}{2\times (-1)}=45\)

OpenStudy (anonymous):

oops i wrote "minimum" i meant "maximum" sorry

OpenStudy (anonymous):

45 is the maximum revenue right? so what would be the answer to the other question?

OpenStudy (anonymous):

??